OpenStudy (anonymous):

Write an equation that has an x-intercept of 3 and is perpendicular to the graph of 5x - 3y = 8. (show work)

OpenStudy (anonymous):

First, solve for y in the original equation. We get y=(5/3)x-(8/3). Then, for a line to be perpendicular to a given line, the slope must be the opposite reciprical of the other slope (in other words, the original slope is 5/3, so the slope of the perpendicular line is -3/5). So we get y+b=(-3/5)x I hope you can do the rest =P